OpenStudy (anonymous):

Evaluate \[\lim_{x \rightarrow 0} \frac{(1+x)^{\frac{1}{k}} -1}{x}\]

hartnn (hartnn):

L'Hopital's rule not allowed ?

OpenStudy (anonymous):

Yes, not allowed!

hartnn (hartnn):

put 1+x = y

hartnn (hartnn):

then use \[\lim_{y \rightarrow a}\frac{ y^n-a^n }{ y-a }=ny^{n-1}\]

OpenStudy (anonymous):

Can you explain how you get that?

hartnn (hartnn):

the formula ? its standard, like sin x/x=1 after substittution, u get lim y->1 y^(1/k)-1/(y-1)

hartnn (hartnn):

which is just (1/k) y ^{1/k-1} which you would expect by L'hopitals....

OpenStudy (anonymous):

If you don't mind, I would like to know how you get that formula, since I haven't learnt that before, if I remember correctly.

hartnn (hartnn):

sure, u know the general formula for a^n-b^n ??

hartnn (hartnn):

\(a^n-b^n = (a-b)(a^{n-1}+a^{n-2}b+a^{n-3}b^2+....b^{n-1})\) use this : y-a will cancel out pt y=a in remaining polynomial.

hartnn (hartnn):

tell me if u couldn't get it.

OpenStudy (anonymous):

Yes!! I was trying to use that formula but seemed that I couldn't continue.. Since a = 1+x there, b =1 and n= 1/k When I substitute a = 1+x, it seems that I have to ''expand'' (1+x)^(1/k -1). That would be a mess!

hartnn (hartnn):

oh, yes. that can also be done..... but y^n-a^n/y-a is quite standard.....

OpenStudy (anonymous):

But it comes to a problem that I haven't learnt it.. And it's hard for me to write that if I don't understand that.. I'm sorry!!!

hartnn (hartnn):

thats actually good spirit, keep it up.

OpenStudy (anonymous):

Haha! Would you mind explaining the standard form??

hartnn (hartnn):

but u need to know this relation: \(a^n-b^n = (a-b)(a^{n-1}+a^{n-2}b+a^{n-3}b^2+....b^{n-1})\)

OpenStudy (anonymous):

Yes, I know it!

hartnn (hartnn):

a-b cancelled out right ?

hartnn (hartnn):

1+x-1 = x

OpenStudy (anonymous):

Sure!!

hartnn (hartnn):

ok, so now directly put x=0, what u get ?

hartnn (hartnn):

i am sorry, the formula is \(\lim_{y \rightarrow a}\frac{ y^n-a^n }{ y-a }=na^{n-1}\)

OpenStudy (anonymous):

1+1+1+...+1 ....

hartnn (hartnn):

1/k times, right ?

OpenStudy (anonymous):

Does that matter?

OpenStudy (anonymous):

I suppose 1^(1/k) = 1 ..

OpenStudy (anonymous):

For k =/= 0

hartnn (hartnn):

yup, u need to know how many times '1' is added.

hartnn (hartnn):

1^{1/k-1} + 1^{1/k-2}+....1^{1}

OpenStudy (anonymous):

Ahh.. Yes.... 1/k times..

hartnn (hartnn):

then it would be 1/k, isn't it ?

OpenStudy (anonymous):

It is!

hartnn (hartnn):

sorry for my typing mistake earlier....its na^{n-1} in that formula.... any more doubts here ?

OpenStudy (anonymous):

I don't understand...

hartnn (hartnn):

u got 1/k, its correct answer, which part did u not understand ?

OpenStudy (anonymous):

na^{n-1} ...

hartnn (hartnn):

that general formula, if u didn't understand, leave it...you were able to solve without that also, isn't it ?

OpenStudy (anonymous):

Oh, you're right! Thanks!!~

hartnn (hartnn):

welcome ^_^